OpenStudy (rose4825):

Can someone help me with this problem: Which is the product of -√2/3 + i and its conjugate? A. 13/9 B. 11/9 C. 5/3 D. -1/3 I don't understand how to do it.

OpenStudy (rose4825):

@jhonyy9 @sshayer

OpenStudy (rose4825):

Negative?

OpenStudy (irishboy123):

\((a+ib)(a-ib) = a^2 + b^2\)

OpenStudy (rose4825):

Would it be -1/3?

OpenStudy (sshayer):

is your problem \[-\frac{ \sqrt{2} }{ 3 }+\iota ~or~\frac{ -\sqrt{2} }{ 3+\iota }\]

OpenStudy (irishboy123):

you'll get \((-\frac{\sqrt{2}}{3})^2 + (1)^2\) \(= \frac{{2}}{9} + 1\)

OpenStudy (rose4825):

@sshayer it's \[-\frac{ \sqrt{2} }{ 3 } + i\]

OpenStudy (sshayer):

from the options it seems Irish boy's work is excellent.

OpenStudy (rose4825):

I still don't understand at all

OpenStudy (sshayer):

conjugate of \[a+ \iota b ~is~a- \iota b\] \[\left( a+ \iota b \right)\left( a- \iota b \right)=a^2-b^2\]

OpenStudy (rose4825):

Okay, yeah that makes sense

OpenStudy (sshayer):

\[conjugate~of ~\frac{ -\sqrt{2} }{ 3 }+\iota~is~\frac{ -\sqrt{2} }{ 3 }-\iota \]

OpenStudy (rose4825):

Alright, so what Irishboy said, was that then I would square everything and get 2/9+1

OpenStudy (rose4825):

But that's not an option

OpenStudy (sshayer):

\[\left( \frac{ -\sqrt{2} }{ 3 }+\iota \right)\left( -\frac{ \sqrt{2} }{ 3 }-\iota \right)=\left( -\frac{ \sqrt{2} }{ 3 } \right)^2-\iota^2\] \[=\frac{ 2 }{ 9 }+1=\frac{ 2+9 }{ 9 }=\frac{ 11 }{ 9 }\]

OpenStudy (sshayer):

you got it.

OpenStudy (sshayer):

\[\iota^2=-1\] -(-1)=+1

OpenStudy (rose4825):

That helped so much. Thanks!!

OpenStudy (sshayer):

yw